# Partitioned Quantum Subspace Expansion

### v2

@@ Abstract
We present an iterative generalisation of the quantum subspace expansion algorithm used with a Krylov basis. The iterative construction connects a sequence of subspaces via their lowest energy states. Diagonalising a Hamiltonian in a given Krylov subspace requires the same quantum resources in both the single step and sequential cases. We propose a variance-based criterion for determining a good iterative sequence and provide numerical evidence that these good sequences display improved numerical stability over a single step in the presence of finite sampling noise. Implementing the generalisation requires additional classical processing with a polynomial overhead in the subspace dimension. By exchanging quantum circuit depth for additional measurements the quantum subspace expansion algorithm appears to be an approach suited to near term or early error-corrected quantum hardware. Our work suggests that the numerical instability limiting the accuracy of this approach can be substantially alleviated beyond the current state of the art.
@@ 1 Introduction
In the era of noisy quantum computing without error correction, hybrid quantum-classical algorithms have been proposed for a range of applications, for example: efficiently finding the ground state energy of non-linear mean-field models [1, 2]; dynamical mean field theory algorithms for strongly correlated systems [3, 4]; and optimization problems from adjacent fields [5, 6, 7, 8, 9, 10, 11]. Two general approaches underlying several of these algorithms are the Variational Quantum Eigensolver (VQE) [12] and Quantum Subspace Expansion (QSE) [13, 14, 15] methods. QSE is a quantum-classical hybrid algorithm like VQE, but the classical optimisation problem is mapped to a generalised eigenvalue problem, which has various practical advantages compared with the highly non-linear optimisation required in VQE. In addition, it has been shown [16] that QSE can be combined with powerful error mitigation techniques [17, 18, 19], which makes this approach very attractive for noisy quantum devices.
@@
The crucial task in QSE consists of generating a subspace with a lowest energy state that has properties which are sufficiently close to those of the true ground state. Most schemes for generating this subspace aim to create a Krylov subspace due to the excellent performance of this approach in classical algorithms [20]. A Krylov subspace is defined by a set of non-orthogonal basis states constructed through application of powers of the Hamiltonian to a reference state. An approximate Krylov subspace can be generated by performing real [13, 21, 22] and imaginary [23] time evolution, by forming a superposition of real time evolutions [24] (which can remove the need for the Hadamard test [25]), by decomposing Hamiltonian powers into products of unitary operators to form a fine grained Krylov subspace [26], or by generating Gaussian powers of the Hamiltonian [27]. Recently, it has been shown [28] that QSE with an exact Krylov basis can be realized via qubitization [29] and applied to problems from condensed matter physics, quantum chemistry and quantum field theory [28, 30].
@@
An advantageous feature of QSE with a Krylov basis is that the accuracy of the ground state energy increases exponentially with the number of basis states [20, 31]. However, in practice the basis size cannot be increased arbitrarily because higher-order Krylov states become almost linearly dependent due to properties of the power iteration. Consequently, the overlap and Hamiltonian matrices become ill-conditioned such that noise in the matrix elements measured on quantum hardware leads to numerical instability. It has been recently shown that this problem can be mitigated by projecting the generalised eigenvalue problem onto a subspace with dominant singular values [32]. This thresholded QSE (TQSE) approach results in numerical stability up to a critical basis length, but increasing the accuracy by extending the basis size remains a desirable yet elusive capability.
@@
Here we show that the accuracy of QSE with statistical noise can be increased beyond that achieved by TQSE while maintaining numerical stability. We achieve this by introducing a Partitioned Quantum Subspace Expansion (PQSE), where one large QSE instance is partitioned into several smaller QSE problems which exhibit better conditioning. The introduced PQSE scheme is iterative, with the output of one small QSE instance serving as input to the next small QSE problem. The scheme is illustrated in Fig. (1) for the simple case of a subspace being partitioned into two smaller subspaces [a technical introduction to the figure is provided in Sec. (2.2)]. PQSE is intended for operation in the Krylov basis regime where numerical instability becomes an obstacle to the effectiveness of TQSE. Overall PQSE requires the same quantum resources but more classical calculations than QSE or TQSE. The classical cost is determined by the choice of partitioning. By leveraging additional quantum resources, we provide a criterion for choosing an effective partitioning based on the energy variance. We analyse two canonical examples, a spin ring with random potential and a molecular Hamiltonian. We find that the size of the Krylov basis can be substantially increased compared to standard QSE and TQSE, and thus PQSE achieves substantially higher accuracy.
@@ 2.1 Background
A special choice of subspace in Eq. (1) is given by a Krylov subspace. A Krylov subspace of order $R$ starting from a reference state |$\ket{\phi_{0}}$ is defined as
$\mathcal{K}_{R}(H,\ket{\phi_{0}})=\text{span}\{\ket{\phi_{0}},H\ket{\phi_{0}},\ldots,H^{R-1}\ket{\phi_{0}}\}.$ (7)
@@
It has been previously noted that the QSE matrices with a Krylov basis have a Hermitian-Toeplitz structure [32]. This structure allows both QSE matrices to be constructed by calculating the first row and final column of $\mathbf{H}$.
@@ 2.2 Partitioned Quantum Subspace Expansion
We illustrate the idea of our algorithm by considering $P$ sequential QSE problems. A guiding schematic for $P=2$ is shown in Fig. 1. Starting from an initial state |$\ket{\phi_{0}}$, we build the subspace
@@
and denote its corresponding ground state by |$\ket{\psi_{2}}$. More generally, we consider the sequence
@@
where we define the order of the partition $\mathcal{R}$ as  [^1] It will become clear in Sec. (2.2.1) that $O(\mathcal{R})$ need not equal $R$.
@@
Consequently, each QSE problem in $\mathcal{K}_{r_{k}}(H,\ket{\psi_{k-1}})$ can be solved with the matrix elements in Eq. (2.1) required for building the subspace $\mathcal{K}_{R}(H,\ket{\phi_{0}})$. However, the ground state in $\mathcal{K}_{R}(H,\ket{\phi_{0}})$ will in general be different from the final ground state |$\ket{\psi_{P}}$ obtained by PQSE. Similarly, the ground state energies of the two approaches will be different as well. Without noise and with infinite numerical precision, the direct solution obtained by solving the QSE problem in $\mathcal{K}_{R}(H,\ket{\phi_{0}})$ is expected to be better. However, in the presence of noise and finite numerical precision this need not be the case. The advantage of PQSE is that one has to solve a sequence of smaller, better conditioned, QSE problems.
@@ 2.2.1 PQSE Algorithm definition
This algorithm returns an estimate $E_{g}$ of the ground state energy of a Hamiltonian $H$, as well as a description of the approximate ground state. For a given Krylov subspace order $R$, it finds a partition into smaller QSE problems $\mathcal{R}$ [see Eq. (11)] such that all small Krylov subspaces are part of $\mathcal{K}_{R}(H,\ket{\phi_{0}})$. The PQSE algorithm is as follows.
@@
    Input: Hamiltonian $H$, Krylov subspace of order $R$ with initial reference state |$\ket{\phi_{0}}$. The matrix elements defined in Eq. (2.1) are obtained using a quantum computer.
@@
        Create the generalised eigenvalue problem in $\mathcal{K}_{q}(H,\ket{\psi_{P-1}})$ [see Appendix B].
@@
        Calculate the variance $\text{var}(H,\ket{q})$ [see Appendix B.2].
@@ 2. $\bullet$
We now quantify the resources used in PQSE. PQSE uses all matrix elements of the full Krylov subspace of order $R$. In addition, calculating the variance may require the additional matrix element $\bra{\phi_{0}}H^{2R}\ket{\phi_{0}}$, which is not needed in standard QSE, which requires the elements specified in Eq. (2.1). All QSE techniques incur the classical cost of constructing and solving Eq. (3). The additional classical cost of PQSE comes from combining the output of previous QSE problems with quantum expectation values to produce new QSE problems. While this cost may vary based on noise levels and problem parameters, we find that the classical cost scales as $\mathcal{O}(R^{4})$, see Appendix B. QSE and TQSE incur an $\mathcal{O}(R^{3})$ classical cost to solve the generalised eigenvalue problem of Eq. (3) [33].
@@ 2.3 Noise model
We now describe how we model noise on each matrix element in Eq. (2.1). Noisy versions of $\mathbf{H}$ and $\mathbf{S}$ are constructed as
@@
where $\mathbf{\Delta}_{H}$ and $\mathbf{\Delta}_{S}$ are Hermitian-Toeplitz matrices with elements $\Delta_{H,ij}$ and $\Delta_{S,ij}$ which quantify the uncertainty in $\bra{\phi_{0}}H^{i+j+1}\ket{\phi_{0}}$ and $\bra{\phi_{0}}H^{i+j}\ket{\phi_{0}}$. Because $\mathbf{H}$ contains all elements of $\mathbf{S}$ except $S_{00}=\braket{\phi_{0}\lvert\phi_{0}}=1$, we only consider constructing $\mathbf{\Delta}_{H}$. We draw $\mathbf{\Delta}_{H,ij}$ from a normal distribution of width $\sigma_{H,ij}$ where
@@
for a dimensionless parameter $\delta$ which controls noise strength. By considering measuring each matrix element $1/\delta^{2}$ times we see that this choice of noise model is equivalent to assuming we have access to unbiased estimators of $H_{ij}$, with any error resulting from finite sampling or limited numerical precision. This assumption can be motivated by considering applying this technique in either i) the regime where quantum error mitigation can still be effectively applied, but classical simulation is not possible [34]; ii) the regime of fault tolerant quantum computing. In [28] it is suggested that Krylov basis QSE may be a favourable choice over Phase Estimation in the early fault-tolerant regime, highlighting the former’s noise robustness as motivation.
When implementing TQSE to mitigate the effects of noise, we follow [32] and set our thresholding parameter $\tau$ to $\mathcal{O}\left((\eta^{2}_{H}+\eta^{2}_{S})^{1/2}\right)$ where $\eta_{H},\eta_{S}$ are the spectral norms of the perturbations $\mathbf{\Delta}_{H},\mathbf{\Delta}_{S}$.
@@ 3 Results
[…]
@@
[…]
We begin with a comparison of the different QSE techniques in the absence of noise ($\delta=0$). For each Hamiltonian we vary the Krylov basis order $R$ and compute the relative error in the ground state energy
@@
where $\tilde{E}_{g}$ is the ground state energy obtained by exact diagonalization and $E_{g}$ is obtained by either QSE, TQSE or PQSE.
The results for $\epsilon_{\text{rel}}$ as a function of $R$ in case of the spin ring are shown in Fig. 2(b). We find that the relative error initially decreases exponentially with $R$ for all three techniques. In spite of the absence of intentionally added noise, finite machine precision introduces a barrier to arbitrarily low solution error and results in numerical instability. In order to mitigate the effect of noise due to finite machine precision we fix the TQSE threshold parameter to $\tau=10^{-13}$, a value found to be effective in other work [28]. Note that thresholding is not used as part of PQSE.
With this, TQSE allows one to explore larger values of $R$ and the relative error is approximately one order of magnitude smaller compared to QSE. PQSE can have a slighter higher average error at basis orders where QSE fails and TQSE is numerically stable. On the other hand, PQSE returns more accurate solutions in regions where the other techniques fail. We find that by increasing $R$, PQSE can lower $\epsilon_{\text{rel}}$ by two orders of magnitude compared to TQSE.
The results for $\epsilon_{\text{rel}}$ as a function of $R$ in case of the $\text{H}_{6}$ molecule are shown in Fig. 2(e). We observe that $\epsilon_{\text{rel}}$ decays much slower with $R$ as compared to the spin ring for all QSE techniques. This finding is consistent with the results in [28], where the slow convergence with $R$ was explained by the use of a Hartree-Fock state as a reference state. QSE and TQSE display similar qualitative behaviour to the spin ring example. In contrast to the spin ring example, PQSE does not display signs of numerical instability, instead converging to a fixed value of $\epsilon_{\text{rel}}$. This fixed value is higher than the minimum error achieved by QSE and TQSE. We obtain insight into this by computing the overlap of each QSE solution with the true ground state found using exact diagonalisation. At $R=6$, QSE returns a lower energy error than PQSE, but the QSE ground state estimate has a smaller overlap with the true ground state than the PQSE estimate due to numerical instabilities. On the other hand, the overlap of the ground state estimate obtained by TQSE for $R=6$ and $R=7$ is larger than the corresponding values for PQSE, which is consistent with the better energy estimates of TQSE at these points.
Next we investigate the performance of the different QSE techniques at a fixed noise strength $\delta=10^{-6}$. Our performance metric is denoted by $\overline{\epsilon_{\text{rel}}}$, where the overbar denotes the average of $\epsilon_{\text{rel}}$ in Eq. (18) over noise instances. The corresponding standard error for $\overline{\epsilon_{\text{rel}}}$ is indicated by error bars in Figs. 2(c) and (f). In contrast to the $\delta=0$ example, we see that in both examples numerical instability restricts use of QSE to Krylov basis orders of 1 or 2. Implementing thresholding allows an exponentially fast decrease in average error for a limited range of $R$. If $R$ is increased beyond values where thresholding also exhibits instability, PQSE grants useful access to larger Krylov basis sizes and more accurate energy estimates. For the spin ring and molecular examples, in Figs. 2(c) and (f) respectively, this corresponds to a two order and half an order of magnitude decrease in $\overline{\epsilon_{\text{rel}}}$. As in the noiseless case, we find that QSE methods are more effective for the spin ring example than for $\text{H}_{6}$. Additional results for a range of noise strengths are provided in Appendix A and show similar qualitative behaviour of the relative energy error as for $\delta=10^{-6}$.
In order to investigate the performance of each approach to QSE as a function of noise strength, we compute $\overline{\epsilon_{\text{rel}}}$ for a fixed noise strength and find its minimum value as a function of $R$,
@@
The results for $\xi$ as a function of noise strength are shown in Figs. 3(a) and (b) for the spin ring and the $\text{H}_{6}$ molecule, respectively. We find that the the ability of PQSE to achieve a lower error on average is reflected across the range of noise strengths we consider. The reduction in error provided by PQSE is the greatest in the spin ring example, where $\xi$ can approximately two orders of magnitude below both QSE and TQSE, see Fig. 3(a). The results for $\text{H}_{6}$ in Figure 3(b) show that PQSE also provides an advantage for varying noise strengths, achieving up to an order of magnitude increase in precision.
@@
In order to both better understand the spin ring example and the behaviour of PQSE with different choices of reference state, we apply each QSE algorithm to solving Eq. (16) for a range of values of $J/h$. We use the $J=0$ solution as an initial state |$\ket{\phi_{0}}$ for $J/h=0.1,0.2$ and the $h=0$ solution for larger ratios. The results for $\xi$ are shown in Fig. 3(c) where, for clarity, we include two noise strengths $\delta=10^{-6},10^{-1}$. We see that PQSE provides a lower minimum average error than QSE and TQSE for nearly all values of $\delta$ and all $J/h$. For $\delta=10^{-1}$ QSE has substantially larger $\xi$ than other techniques for all $J/h$. The difference between PQSE and TQSE ranges from a half order to just over one order of magnitude. For $\delta=10^{-6}$ the difference in $\xi$ between PQSE and TQSE ranges from around two orders of magnitude at $J/h=0.1$, reducing from one and a half orders to just under a half order from $J/h=0.2\dots 2$ before reaching zero difference at $J/h=5$. At $J/h=10$ TQSE outperforms PQSE by half an order of magnitude. All techniques display greater variation in $\xi$ across the range of $J/h$ for weaker noise compared with the values of $\xi$ for stronger noise. This suggests that for weaker noise the numerical stability of each QSE technique, particularly PQSE, is more sensitive to the properties of $H$ and the choice of initial state |$\ket{\phi_{0}}$, than for stronger noise.
[…]

### v3

@@ Abstract
We present an iterative generalisation of the quantum subspace expansion algorithm used with a Krylov basis. The iterative construction connects a sequence of subspaces via their lowest energy states. Diagonalising a Hamiltonian in a given Krylov subspace requires the same quantum resources in both the single step and sequential cases. We propose a variance-based criterion for determining a good iterative sequence and provide numerical evidence that these good sequences display improved numerical stability over a single step in the presence of finite sampling noise. Implementing the generalisation requires additional classical processing with a polynomial overhead in the subspace dimension. By exchanging quantum circuit depth for additional measurements the quantum subspace expansion algorithm appears to be an approach suited to near term or early error-corrected quantum hardware. Our work suggests that the numerical instability limiting the accuracy of this approach can be substantially alleviated in a parameter-free way.
@@ 1 Introduction
In the era of noisy quantum computing without error correction, hybrid quantum-classical algorithms have been proposed for a range of applications, for example: efficiently finding the ground state energy of non-linear mean-field models [1, 2]; dynamical mean field theory algorithms for strongly correlated systems [3, 4]; and optimization problems from adjacent fields [5, 6, 7, 8, 9, 10, 11]. Two general approaches underlying several of these algorithms are the Variational Quantum Eigensolver (VQE) [12] and Quantum Subspace Expansion (QSE) [13, 14, 15] methods. QSE is a quantum-classical hybrid algorithm like VQE, but the classical optimisation problem is mapped to a generalised eigenvalue problem, which has various practical advantages compared with the highly non-linear optimisation required in VQE. In addition, it has been shown [16, 17, 18] that QSE can be combined with powerful error mitigation techniques [19, 20, 21], which makes this approach very attractive for noisy quantum devices.
@@
The crucial task in QSE consists of generating a subspace with a lowest energy state that has properties which are sufficiently close to those of the true ground state. Most schemes for generating this subspace aim to create a Krylov subspace due to the excellent performance of this approach in classical algorithms [22]. A Krylov subspace is defined by a set of non-orthogonal basis states constructed through application of powers of the Hamiltonian to a reference state. A Krylov subspace can also be generated by performing real [13, 23, 24] and imaginary [25] time evolution, by forming a superposition of real time evolutions [26] (which can remove the need for the Hadamard test [27]), by decomposing Hamiltonian powers into products of unitary operators to form a fine grained Krylov subspace [28], or by generating Gaussian powers of the Hamiltonian [29]. Recently, it has been shown [30] that QSE with an exact Krylov basis can be realized via qubitization [31] and applied to problems from condensed matter physics, quantum chemistry and quantum field theory [30, 32].
@@
An advantageous feature of QSE with a Krylov basis is that the accuracy of the ground state energy increases exponentially with the number of basis states [22, 33]. However, in practice the basis size cannot be increased arbitrarily because higher-order Krylov states become almost linearly dependent due to properties of the power iteration. Consequently, the overlap and Hamiltonian matrices become ill-conditioned such that noise in the matrix elements measured on quantum hardware leads to numerical instability. It has been recently shown that this problem can be mitigated by projecting the generalised eigenvalue problem onto a subspace with dominant singular values [34]. This thresholded QSE (TQSE) approach results in numerical stability up to a critical basis length, but increasing the accuracy by extending the basis size remains a desirable yet elusive capability. Furthermore, the effectiveness of TQSE is reliant on finding a suitable threshold parameter. This parameter will depend on properties of the target problem and noise. Although good parameters are known to exist [34, 30, 35], finding these can be challenging in practice.
@@
Here we show that the accuracy of QSE with statistical noise can be made comparable to that achieved by TQSE in a parameter-free way. We achieve this by introducing a Partitioned Quantum Subspace Expansion (PQSE), where one large QSE instance is partitioned into several smaller QSE problems which exhibit better conditioning. The introduced PQSE scheme is iterative, with the output of one small QSE instance serving as input to the next small QSE problem. The scheme is illustrated in Fig. (1) for the simple case of a subspace being partitioned into two smaller subspaces [a technical introduction to the figure is provided in Sec. (2.2)]. PQSE is intended for operation in the Krylov basis regime where numerical instability becomes an obstacle to the effectiveness of TQSE. Overall PQSE requires the same quantum resources but more classical calculations than QSE or TQSE with a fixed threshold parameter. The classical cost is determined by the choice of partitioning. By leveraging additional quantum resources, we provide a criterion for choosing an effective partitioning based on the energy variance. We analyse two canonical examples, a spin ring with random potential and a molecular Hamiltonian. We find that the size of the Krylov basis can be substantially increased compared to standard QSE, and thus PQSE achieves substantially higher accuracy, comparable to that achieved by TQSE.
@@ 2.1 Background
A special choice of subspace in Eq. (1) is given by a Krylov subspace. A Krylov subspace of order $R$, generated by an operator $A$ and starting from a reference state |$\ket{\phi_{0}}$ is defined as
$\mathcal{K}_{R}(A,\ket{\phi_{0}})=\text{span}\{\ket{\phi_{0}},A\ket{\phi_{0}},\ldots,A^{R-1}\ket{\phi_{0}}\}.$ (7)
@@
$H_{ij}$ $=\braket{\phi_{0}\lvert(A^{i})^{\dagger}HA^{j}\rvert\phi_{0}}\,,$ (8a)
$S_{ij}$ $=\braket{\phi_{0}\lvert(A^{i})^{\dagger}A^{j}\rvert\phi_{0}}\,,$ (8b)
$i,j$$\in\{0,1,\ldots,R-1\}.$ (8c)
Taking $A=H$ the matrices are
@@
With a Hamiltonian power basis these QSE matrices have a Hankel structure. This structure allows both QSE matrices to be constructed by calculating the first row and final column of $\mathbf{H}$.
@@ 2.2 Partitioned Quantum Subspace Expansion
We illustrate the idea of our algorithm by considering $P$ sequential QSE problems. We will focus on a Hamiltonian power basis i.e. $A=H$, and show how the method can be straightforwardly generalized later in the text. A guiding schematic for $P=2$ is shown in Fig. 1. Starting from an initial state |$\ket{\phi_{0}}$, we build the subspace
@@
and denote its corresponding ground state by |$\ket{\psi_{2}}$. The largest power of $H$ in |$\ket{\psi_{2}}$ will be $r_{1}+r_{2}-2$. More generally, we consider the sequence
@@
where we define the order of the partition $\mathcal{R}$ as
@@
Consequently, each QSE problem in $\mathcal{K}_{r_{k}}(H,\ket{\psi_{k-1}})$ can be solved with the matrix elements in Eq. (9) required for building the subspace $\mathcal{K}_{R}(H,\ket{\phi_{0}})$. However, the ground state in $\mathcal{K}_{R}(H,\ket{\phi_{0}})$ will in general be different from the final ground state |$\ket{\psi_{P}}$ obtained by PQSE. Similarly, the ground state energies of the two approaches will be different as well. Without noise and with infinite numerical precision, the direct solution obtained by solving the QSE problem in $\mathcal{K}_{R}(H,\ket{\phi_{0}})$ is expected to be better. However, in the presence of noise and finite numerical precision this need not be the case. The advantage of PQSE is that one has to solve a sequence of smaller, better conditioned, QSE problems.
@@ 2.2.1 PQSE Algorithm definition
This algorithm returns an estimate $E_{g}$ of the ground state energy of a Hamiltonian $H$, as well as a description of the approximate ground state. For a given Krylov subspace order $R$ it finds a partition into smaller QSE problems $\mathcal{R}$ [see Eq. (12)] such that all small Krylov subspaces are part of $\mathcal{K}_{R}(H,\ket{\phi_{0}})$. The PQSE algorithm is as follows.
@@
    Input: Hamiltonian $H$, Krylov subspace of order $R$ with initial reference state |$\ket{\phi_{0}}$. The matrix elements defined in Eq. (9) are obtained using a quantum computer.
@@
        Create the generalised eigenvalue problem in $\mathcal{K}_{q}(H,\ket{\psi_{P-1}})$ [see Appendix C.1].
@@
        Calculate the variance $\text{var}(H,\ket{q})$ [see Appendix C.3].
@@ 2. $\bullet$
We now quantify the resources used in PQSE. PQSE uses all matrix elements of the full Krylov subspace of order $R$. In addition, calculating the variance may require the additional matrix element $\bra{\phi_{0}}H^{2R}\ket{\phi_{0}}$.
All QSE techniques incur the classical cost of constructing and solving Eq. (3). The additional classical cost of PQSE comes from combining the output of previous QSE problems with quantum expectation values to produce new QSE problems. While this cost may vary based on noise levels and problem parameters, we find that the classical cost scales as $\mathcal{O}(R^{4})$, see Appendix C.2. QSE and TQSE incur an $\mathcal{O}(R^{3})$ classical cost to solve the generalised eigenvalue problem of Eq. (3) [36].
The generalisation of the algorithm description to a Krylov basis generated by an operator $A\leavevmode\nobreak\ \neq\leavevmode\nobreak\ H$ is straightforward, see Appendix C.3.
@@ 2.3 Noise model
We now describe how we model noise on each matrix element in Eq. (9). Noisy versions of $\mathbf{H}$ and $\mathbf{S}$ are constructed as
@@
where $\mathbf{\Delta}_{H}$ and $\mathbf{\Delta}_{S}$ are Hankel matrices with elements $\Delta_{H,ij}$ and $\Delta_{S,ij}$ which quantify the uncertainty in $\bra{\phi_{0}}H^{i+j+1}\ket{\phi_{0}}$ and $\bra{\phi_{0}}H^{i+j}\ket{\phi_{0}}$. Because $\mathbf{H}$ contains all elements of $\mathbf{S}$ except $S_{00}=\braket{\phi_{0}\lvert\phi_{0}}=1$,
we only consider constructing $\mathbf{\Delta}_{H}$. We draw $\mathbf{\Delta}_{H,ij}$ from a normal distribution of width $\sigma_{H,ij}$ where
@@
for a dimensionless parameter $\delta$ which controls noise strength. By considering measuring each matrix element $1/\delta^{2}$ times we see that this choice of noise model is equivalent to assuming we have access to unbiased estimators of $H_{ij}$, with any error resulting from finite sampling or limited numerical precision. This assumption can be motivated by considering applying this technique in either i) the regime where quantum error mitigation can still be effectively applied, but classical simulation is not possible [37]; ii) the regime of fault tolerant quantum computing. In [30] it is suggested that Krylov basis QSE may be a favourable choice over Phase Estimation in the early fault-tolerant regime, highlighting the former’s noise robustness as motivation. When implementing TQSE to mitigate the effects of noise, we follow [34] and set our thresholding parameter $\tau$ to $10^{-a}\left((\eta^{2}_{H}+\eta^{2}_{S})^{1/2}\right)$ where $\eta_{H},\eta_{S}$ are the spectral norms of the perturbations $\mathbf{\Delta}_{H},\mathbf{\Delta}_{S}$ and $a$ is a threshold scaling hyper-parameter we tune to approximate the optimal value of $\tau$. In order to tune this hyper-parameter we run each TQSE experiment with a given $\delta$ for 50 evenly spaced values of $-0.5\leq a\leq 5$ and take the value which gives the best solution according to the particular error metric we are using. This approach uses the noise-free QSE matrices to calculate $\eta_{H}$, $\eta_{S}$ and the true ground state energy to minimize $a$ with respect to our error metric. We highlight that finding the optimal threshold in this way would not be possible without knowledge of the exact solution.
@@ 3 Results
The second example system is given by an electronic structure Hamiltonian for the $\text{H}_{6}$ chain, see Fig. 2(d) for a schematic and Fig. 2(e, f) for results. The electronic structure Hamiltonian for H6 can be written as
@@
Figure 2: Average relative error in estimate of ground state energy for QSE with increasing Krylov basis of order $R$ generated using Hamiltonian powers: unmodified (QSE), thresholded (TQSE), partitioned (PQSE). Left column: 10 qubit periodic disordered Heisenberg model with $J$ = 0.1, $h$ = 1, the model is depicted in (a) with results for Gaussian noise strength $\delta=0$ in (b) and $\delta=10^{-6}$ in (c). Right column: H6 in a STO-3G basis with a parity mapping applied in equilibrium configuration, the system is depicted in (d) where H and $d$ represent Hydrogen molecules and bond length respectively, with results for Gaussian noise strength $\delta=0$ in (e) and $\delta=10^{-6}$ in (f).
We now present the three different error metrics we will use. For comparisons of the different QSE techniques in the absence of noise ($\delta=0$), we vary the Krylov basis order $R$ and compute the relative error in the ground state energy
@@
where $\tilde{E}_{g}$ is the ground state energy obtained by exact diagonalization and $E_{g}$ is obtained by either QSE, TQSE or PQSE. For comparisons at a non-zero noise strength $\delta$ our performance metric is denoted by $\overline{\epsilon_{\text{rel}}}$, where the overbar denotes the average of $\epsilon_{\text{rel}}$ in Eq. (19) over 200 noise instances. The associated uncertainty in $\overline{\epsilon_{\text{rel}}}$ is the standard error. In order to investigate the performance of each approach to QSE as a function of noise strength, we compute $\overline{\epsilon_{\text{rel}}}$ for a fixed noise strength, again over 200 noise instances, and find its minimum value as a function of $R$,
@@
Heisenberg model. The results for $\epsilon_{\text{rel}}$ as a function of $R$ in case of the spin ring are shown in Fig. 2(b). We find that the relative error initially decreases exponentially with $R$ for all three techniques. In spite of the absence of intentionally added noise, finite machine precision introduces a barrier to arbitrarily low solution error and results in numerical instability. In order to mitigate the effect of noise due to finite machine precision we fix the TQSE threshold parameter to $\tau=10^{-13}$, a value found to be effective in other work [30]. Note that thresholding is not used as part of PQSE. With this, TQSE allows one to explore larger values of $R$ and the relative error is approximately one order of magnitude smaller compared to QSE. PQSE can have a slighter higher average error at basis orders where QSE fails and TQSE is numerically stable. On the other hand, PQSE returns more accurate solutions in regions where the other techniques fail. We find that by increasing $R$, PQSE can lower $\epsilon_{\text{rel}}$ by three orders of magnitude compared to TQSE.
We now move on to results with a fixed noise strength $\delta=10^{-6}$, shown in Fig. 2(c). In contrast to the $\delta=0$ example, we see that numerical instability restricts use of QSE to Krylov basis orders of 1 or 2. Implementing thresholding allows an exponentially fast decrease in average error for a limited range of $R$. For $8\leq R\leq 11$ we see that the average error of TQSE increases before reconverging. The average error of PQSE decreases to a minimum at $R=23$ which is marginally lower than that of TQSE. We provide insight into the subspace dimensions being discarded by TQSE, as well as the typical sequences chosen by PQSE in this experiment in Appendix D.
The results for $\xi$ as a function of noise strength are shown in Figs. 3(a). We find that PQSE performs competitively with TQSE across the range of noise strengths we consider, with $\xi$ between half an order and one and a half orders of magnitude lower for TQSE than PQSE for $10^{-10}\leq\delta\leq 10^{-8}$. The two approaches perform roughly the same at $10^{-7}\leq\delta\leq 10^{-5}$. At $10^{-4}\leq\delta\leq 10^{-1}$, $\xi$ for PQSE is approximately an order of magnitude lower than TQSE.
@@
In order to both better understand the spin ring example and the behaviour of PQSE with different choices of reference state, we apply each QSE algorithm to solving Eq. (17) for a range of values of $J/h$. We use the $J=0$ solution as an initial state |$\ket{\phi_{0}}$ for $J/h=0.1,0.2$ and the $h=0$ solution for larger ratios. The results for $\xi$ are shown in Fig. 3(c) where, for clarity, we include two noise strengths $\delta=10^{-6},10^{-1}$. We see that at $\delta=10^{-1}$, $\xi$ for PQSE is between half an order to an order of magnitude lower than TQSE. At $\delta=10^{-6}$, $\xi$ for TQSE is lower than that for PQSE for $J/h>10^{-1}$ from half an order to two orders of magnitude. All techniques display greater variation in $\xi$ across the range of $J/h$ for weaker noise compared with the values of $\xi$ for stronger noise. This suggests that for weaker noise the numerical stability of each QSE technique, particularly PQSE, is more sensitive to the properties of $H$ and the choice of initial state |$\ket{\phi_{0}}$, than for stronger noise.
Hydrogen chain. The results for $\epsilon_{\text{rel}}$ as a function of $R$ in case of the $\text{H}_{6}$ chain are shown in Fig. 2(e). We observe that $\epsilon_{\text{rel}}$ decays much slower with $R$ as compared to the spin ring for all QSE techniques. This finding is consistent with the results in [30], where the slow convergence with $R$ was explained by the use of a Hartree-Fock state as a reference state. QSE and TQSE display similar qualitative behaviour to the spin ring example. In contrast to the spin ring example, PQSE does not display signs of numerical instability, instead converging to a fixed value of $\epsilon_{\text{rel}}$. This fixed value is higher than the minimum error achieved by QSE and TQSE. We obtain insight into this by computing the overlap of each QSE solution with the true ground state found using exact diagonalisation. At $R=6$, QSE returns a lower energy error than PQSE, but the QSE ground state estimate has a smaller overlap with the true ground state than the PQSE estimate due to numerical instabilities leading to an inconsistency between $E$ and $\bra{\psi}H\ket{\psi}$ after solving Eq. (3). On the other hand, the overlap of the ground state estimate obtained by TQSE for $R=6$ and $R=7$ is larger than the corresponding values for PQSE, which is consistent with the better energy estimates of TQSE at these points.
Results for a fixed noise strength $\delta=10^{-6}$ are shown in Fig. 2(f). We see a monotonic decrease in TQSE error to a minimum at $\overline{\epsilon_{\text{rel}}}=2\times 10^{-4}$, followed by a gradual increase. Here PQSE achieves a minimum error of approximately $\overline{\epsilon_{\text{rel}}}=7\times 10^{-4}$. The increase in $\overline{\epsilon_{\text{rel}}}$ and its variance for PQSE at $R=13$ is due to an ill-conditioned eigenvalue problem. As in the noiseless case, we find that QSE methods are more effective for the spin ring example than for $\text{H}_{6}$. The results for $\xi$ as a function of noise strength are shown in Figs. 3(b). PQSE slightly outperforms TQSE for $10^{-2}\leq\delta\leq 10^{-1}$ and TQSE outperforms PQSE for $10^{-10}\leq\delta\leq 10^{-3}$ by at most approximately half an order of magnitude.
Figure 3: Minimum average relative error $\xi$ in the estimate of ground state energy for QSE with a Krylov basis generated using Hamiltonian powers: unmodified (QSE), thresholded (TQSE), partitioned (PQSE). (a) $\xi$ as a function of noise strength $\delta$ for a 10 qubit periodic disordered Heisenberg model with $J/h=0.1$. (b) $\xi$ as a function of noise strength $\delta$ for H6 in a STO-3G basis with a parity mapping applied in equilibrium configuration. (c) $\xi$ as a function of $J/h$ for the 10 qubit periodic disordered Heisenberg model and noise strengths $\delta=10^{-6},10^{-1}$. Error bars show the standard error.